OpenStudy (anonymous):

6 7 1 - = -- - -- y y 11 solve

OpenStudy (anonymous):

\[\frac{6}{y}=\frac{7}{y}-\frac{1}{11}?\]

OpenStudy (anonymous):

y = 11

OpenStudy (anonymous):

deniminator is 11y\[6*11=(7*11)-(1*y)\]\[66=77-y\]\[y=\frac{-77+66}{-1}=11\]

OpenStudy (anonymous):

That's one way to do it...but there's a much simpler way, lol

OpenStudy (anonymous):

subtract 6/y from both sides, then add 1/11 to both sides. Cross multiply afterwards...